OpenStudy (anonymous):

f(x) = x+x^(1/2) find the derivative using the definition of derivative

OpenStudy (amistre64):

whatcha got so far?

OpenStudy (anonymous):

ummm i know that it's lim of h approaching zero of the equation f(x+h) - f(x) all over h

OpenStudy (amistre64):

well, thats a start :)

OpenStudy (anonymous):

but i cant get the h out of the denominator

OpenStudy (anonymous):

the root is going to be a pain in the neck have fun!

OpenStudy (amistre64):

do you know that the derivative of a sum or difference, is the difference of the derivatives of the parts?

OpenStudy (amistre64):

g(x) = x ; g(x+h) = x+h \[\lim_{h->0}\frac{g(x+h)-g(x)}{h}\] \[\lim_{h->0}\frac{x+h-x}{h}\] \[\lim_{h->0}\frac{h}{h}\] \[\lim_{h->0}\ 1=1\]

OpenStudy (amistre64):

s(x) = sqrt(x) ; s(x+h) = sqrt(x+h) \[\lim_{h->0}\frac{s(x+h)-s(x)}{h}\] \[\lim_{h->0}\frac{\sqrt{x+h}-\sqrt{x}}{h}\] \[\lim_{h->0}\frac{\sqrt{x+h}-\sqrt{x}}{h}*\frac{\sqrt{x+h}+\sqrt{x}}{\sqrt{x+h}+\sqrt{x}}\] \[\lim_{h->0}\ \frac{x+h-x}{h(\sqrt{x+h}+\sqrt{x})}\] \[\lim_{h->0}\ \frac{h}{h(\sqrt{x+h}+\sqrt{x})}\] \[\lim_{h->0}\ \frac{1}{\sqrt{x+h}+\sqrt{x}}\] \[\lim_{h->0}\ \frac{1}{\sqrt{x+0}+\sqrt{x}}=\frac{1}{2\sqrt{x }}\]

OpenStudy (amistre64):

sooo..... out them together now :)

OpenStudy (amistre64):

or put them together .... you can out them if you want, but it just doesnt have the desired effect

OpenStudy (anonymous):

very nice! btw @casey this derivative comes up so frequently you should just remember it. it will save you a ton of time later. like memorizing 8*7